OpenStudy (anonymous):

lim x-> 2 (x^3 - 8)/(2x - 4)

OpenStudy (anonymous):

U need to FACTOR x^3-8

OpenStudy (anonymous):

i already simplified the fraction as: (x^3 - 8)/2(x - 2)

OpenStudy (anonymous):

THATS GOOD @brygiger now, FACTOR x^3-8

OpenStudy (anonymous):

is it (x-2)(x^2 + 2)^2?

OpenStudy (anonymous):

*(x-2)(x^2 + 4)

OpenStudy (anonymous):

NOPE: a^3-b^3 =(a-b)(a^2 +ab +b^2)

OpenStudy (anonymous):

DOES IT HELP?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

so... x^3 - 2^3 = (x-2) (x^2 + 2x + 4)

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

NOW, u can cancel the terms (x-2)

OpenStudy (anonymous):

doing it. but wait isn't (a + b)^2 = (a^2 + 2(ab) + b^2) not (a^2 +ab +b^2)

OpenStudy (anonymous):

x^3-8=x^3-2^3

OpenStudy (anonymous):

answer is 6?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

YEP